Counts Accounting has a beta of 1.55. The tax rate is 40%, and Counts is financed with 45% debt. What is Counts' unlevered beta? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1
Dear Student
Thank you for using Chegg
Please find below the answer
Statement showing Computations
Particulars
Unlevered Beta = Levered Beta / 1 + ((1 - Tax Rate) x Debt/Equity)
Counts' unlevered beta = 1.55/(1+((1-.40)*.45/.55)
Counts' unlevered beta = 1.55/(1+((.60)*.8182)
Counts' unlevered beta = 1.55/(1.4909)
Counts' unlevered beta = 1.04